Find the general solution of an 1D heat equation: ?t (?, ?) = 4?xx (?, ?) with the boundary conditions ?(0, ?) = ?(2, ?) = 0. Note that ?(?, ?) denotes the temperature profile along ? of a uniform rod of length 2.
Answer #1

Given: I-D Heat Equation - 4 Toon (2,t) Te (2,+ Conditions : Boundary To,t) T(2,t) = 0 deryth of uniform read let + u Torse (I dT 4 d²7 2 -d? 7 at dT dt -d²T Х dre dx? 4 dt dt : -X²T de + dx² 4 ons com plemestery feruetrom Equations for above dilferufrom boundary condition م = (رع) من در کوره - ( م ) À was (2x) A+ ہو۔ [A و + و therefore, و: ام ود (+ ,2) - (نن) - A+ e B sBsina - Quals 2 A los an & B sin/72 Turfor, ²t +(at) = e y Al=0 - (@nnit tla,t) Eē a- (2n-1) B sin /(2n-1) (@u-) n=1 2 (fen-